OpenStudy (anonymous):

(Incomplete sol.) Between two no. whose sum is 2 1/6 an even number of arithmetic means is inserted; the sum of these means exceeds their numbers by 1. How many means are there? @nitz

OpenStudy (anonymous):

suppose two numbers are a and b a+b=6

OpenStudy (anonymous):

it is 2+1/6

OpenStudy (anonymous):

13/2

OpenStudy (anonymous):

suppose n means are inserted. and average of all arithmatic means will be same as of whole seriese irrespective of number of a.m. so,average =(a+b)/2=3

OpenStudy (anonymous):

ya 13/2

OpenStudy (anonymous):

s0rry a+b=13/2 and average is 13/4

OpenStudy (anonymous):

"average of all arithmatic means will be same as of whole seriese " , how?

OpenStudy (anonymous):

is it like , in an AP of x,y,z y=(x+z)/2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

wait...i am myself confused

OpenStudy (anonymous):

ok lets try another method let n be the total number of terms in the series and hence n-2 is the no. of means sum of means=n/2(a+1)-(a+1) number of means=n-2

OpenStudy (anonymous):

right till here?

OpenStudy (anonymous):

sum of means , how it came?

OpenStudy (anonymous):

from statement.....read again

OpenStudy (anonymous):

no i am confused.

OpenStudy (anonymous):

write that in fractions

OpenStudy (anonymous):

what in fractions?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

(a+1)-(a+1) how it came.

OpenStudy (anonymous):

it should be a+1+b-1

OpenStudy (anonymous):

so the sum \[\frac{n(a+b)}{2}\]